OpenStudy (anonymous):

Series question for radius of convergence and interval of convergence: Sum (n=1, infinity) 3(n!)x^n

OpenStudy (amistre64):

what do we get as a ratio ?

OpenStudy (anonymous):

(n+1)x, with the limit I got x lim (n->infinity) (n+1) which is infinity

OpenStudy (amistre64):

3(n+1!) x^(n+1) --------------- 3 (n!) x^n (n+1)n! x ---------- (n!) (n+1) x; pull out the x to get: lim of n+1 as n to inf = inf. thats the radius i believe

OpenStudy (amistre64):

err, sum does not converge that is

OpenStudy (anonymous):

Ok, so infinity (-inf, inf) Let me try it

OpenStudy (amistre64):

is this the start of the question? or something that you came to along the way?

OpenStudy (amistre64):

"radius of convergence"; this thing doesnt converge

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} 3(n!)x ^{n}\]

OpenStudy (anonymous):

That is the question at the start, the actual question

OpenStudy (anonymous):

I got that it doesn't converge the first time and it apparently does... somehow

OpenStudy (amistre64):

http://tutorial.math.lamar.edu/Classes/CalcII/PowerSeries.aspx example 3 seems to deal with this case

OpenStudy (amistre64):

the radius of convergence is 0; but it can still have an interval of convergence, most likely just a single point

OpenStudy (amistre64):

hmmm, it says the radius is zero; but the interval are the values of x that make the part pulled out go to zero

OpenStudy (amistre64):

|x| = 0 when x=0

OpenStudy (anonymous):

... Ok, so whatever makes the x(n+1) go to 0?

OpenStudy (anonymous):

Ok, so R = 0 for sure.

OpenStudy (amistre64):

yes, when x(inf) = 0

OpenStudy (amistre64):

if we consider it this way: we found\[|x|(\infty)\]therefore \[|x|(\infty)<1\] \[|x|<\frac{1}{\infty}\] \[|x|<0:~:hence~ R=0\] \[0<x<0:~:interval~=~0\]

OpenStudy (amistre64):

interval is: x=0 that is

OpenStudy (anonymous):

Thank you! That worked out :)

OpenStudy (amistre64):

yay!! :)